OpenStudy (anonymous):

Implicit differentiation 2y^3+y^2-x=0; (3,1)

OpenStudy (anonymous):

6y^2 dy/dx+2y dy/dx=1

OpenStudy (precal):

ok now solve for dy/dx and sub x=3 and y =1 so far your work looks fine

OpenStudy (anonymous):

factor out dy/dx and isolate

OpenStudy (anonymous):

but im a little confused on how to do that

OpenStudy (anonymous):

easiest if you substitute the numbers in now, not later

OpenStudy (precal):

true, unless the instructions are to state dy/dx

OpenStudy (jonnyvonny):

Hay hay I'm here

OpenStudy (anonymous):

\[6y^2 y'+2yy'=1\]and \(y=1\) so \[6y'+2y'=1\] etc makes it much easier to solve

OpenStudy (anonymous):

\[\frac{dy}{dx}(6y^2+2y)=1\Rightarrow \frac{dy}{dx}=\frac{1}{6y^2+2y}\] plug in your point

OpenStudy (jonnyvonny):

y' is the same as dy/dx...when you're at 6y^2'+2yy'=1, you can do this:\[6y^2y'+6yy'=1\rightarrow y'(6y^2+6y)=1\rightarrow y'=1/(6y^2+6y)\]

OpenStudy (jonnyvonny):

1/(6y^2+2y)....sorry

OpenStudy (anonymous):

so \[\frac{dy}{dx}(3,1)= \frac{1}{6(1)^2+2(1)}= \cdots\]

OpenStudy (anonymous):

1/8